OpenStudy (anonymous):

how to solve this using matrices: x1+2x2-2x3+x4=2 2x1-x2+2x3=6 x2-2x4=-1

OpenStudy (anonymous):

\[x _{1}+2x _{2}-2x _{3}+x _{4}=2\] \[2x _{1}-x _{2}+2x _{3}=6\] \[x _{2}-2x _{4}=-1\]

OpenStudy (anonymous):

can be written in augmented matrix form as =1 2 -2 1 2 2 -1 2 0 6 0 1 0 -2 -1 performing r2-->r2-2r1 =1 2 -2 1 2 0 -5 6 -2 2 0 1 0 -2 -1 performing r2-->r2+5r3 =1 2 -2 1 2 0 0 6 -12 -3 0 1 0 -2 -1 performing r2<-->r3 1 2 -2 1 2 0 1 0 -2 -1 0 0 6 -12 -3 since we have three equations but four variables we have to have a free variable.. i assume that variable to be x4 over here so i will get the answer in terms of x4 and constants 1) 6x3-12x4=-3 =>6x3=-3+12x4 =>x3=2x4-(1/2) 2)x2-2x4=-1 =>x2=2x4-1 3)x1+2x2-2x3+x4=2 =>x1=2-2(2x4-1)+2(2x4-1/2)-x4 =>x1 = 2-x4+1 =>x1=3-x4

OpenStudy (anonymous):

thank you! :)